OpenStudy (anonymous):

Describe the following sequence in mathematical terms. 144, 72, 36, 18, 9 A. Descending arithmetic sequence B. Ascending arithmetic sequence C. Descending geometric sequence D. Ascending geometric sequence E. Miscellaneous sequence Can you solve that show me?

OpenStudy (anonymous):

common ratio here is = 72/144=1/2=0.5

OpenStudy (anonymous):

Descending GP

OpenStudy (anonymous):

@Dani do u understand...!!